OpenStudy (adunb8):

Cauchy-Euler Equation I really dont know how to do this anyone care to explain in detail? xy'''+2y"=0

OpenStudy (sirm3d):

if you let \[u=y \prime \prime\] the DE is transformed to \[xu \prime+2u=0\] a DE that can be solved by separation

OpenStudy (adunb8):

i dont quite understand my teacher used like P(D)y = xD^3+2D2 = 0 then multiplied \[x^2\] so it becomes \[x^2D^3+2D^2 = 0 \] then somehow separated into \[xD(D-1)(D-2) + 2D(D-1) = 0\] i dont understand this part

OpenStudy (adunb8):

\[x^3D^3+2x^2D^2\]

OpenStudy (amistre64):

could you solve it if it was: xy' + 2y = 0 ??

OpenStudy (amistre64):

im not familiar with the D as an operator so i cant attest to your teacher method

OpenStudy (adunb8):

hm... would it be x(D+2)y = 0?

OpenStudy (adunb8):

let x = e^t

OpenStudy (adunb8):

im not sure... this is confusing...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=xy%27%27%27%2B2y%27%27%3D0 i keep mismathing it along the way, but this at least gives us something to dbl chk against